Find a general solution to the differential equation. y - y= - 19t + 6 The general solution is y(t) = (Do not use d, D, e,
Answer #1

y-y = -19+ + 6 day - 4 .198 +6 D²y - y d dt dr² is (9?-1)4 = = - 19tto [ whesse DE and D² - d² » (0²1) 4 = -194+6 - - (I) Th1 Do - 1 (1-D²) 1 A particular integroal of (1) is {-19t +63 - 7-197+ (1-02)-1 -194 +67 = -(1+²+...).1 -1964) - -(1+p?) {-191